Fermi National Laboratory


Accelerator Update

Friday September 12 through Monday September 15
All shifts began in shutdown.

On Friday operators worked on the MI-40 LCW pump PLCs and investigated an alarm that uncovered an unplugged "Sump Alarm Chassis." They also found the B1 electronics room temperature at 90 degrees F; operators contacted the Duty Mechanic who restored cooling to the building.

Recycler tasks to be completed during the shutdown:

  • Remove/Reinstall Heater Tape from 13of 17 Gradient magnets
  • Install Heater Tape in 12 of 30 Quadrupole half cells (24 of 60 magnets)
  • Install Cooling tank 10 (horizontal 2-4 GHz) tank
  • Replace approx 170 feet of beam pipe in the 620 Region
  • Bake 20 of 27 vacuum sectors
  • Replace 13 LEP correctors (9 vertical, 4 horizontal)
  • Relocate 2 LEP correctors at the Q305 region
  • Remove IPM from Q404 and Q405
  • Complete cables pulls for: 503-gate valve, IP322

A Tevatron (TeV) expert reported on the first week of shutdown work:

  • All 4 F0 Lambertson liners have been installed and surveyed.
  • Re-installation of the 2nd and 3rd Lambertson will be done on Monday. Vacuum installation work will start Tuesday.
  • The Japanese POTs at TeV sector A-48 have been moved to the Collider Detector Facility (CDF) hall. CDF staff will remove their scintillator detector so that we can bake the entire system out when we reinstall the POTs in 3 weeks.
  • The sector A-4 cryo system is warm.
  • The A side shield wall at CDF has been removed and partially rebuilt. Alignment will be in Saturday to tie in the tunnel and the hall.
  • 20 magnets have had dumb bolts shims installed so far this week. The technicians expect to be done with the D-1 house by Tuesday and then move onto C-4. The shim measurement problem may be due to the changing temperature in the nitrogen shield. Two experts will check the tunnel temps this weekend.
  • All the core drilling for TeV Net monuments are finished. The Dijak bolts will be finished next week. The evening crew is measuring monuments in the TeV E sector. They are picking up speed in this process as they progress. All the towers and scaffolds are in place and ready for the outside network effort.
  • Fixing cold leaks at E-2 will start Monday; all equipment has been staged.
  • Two magnets have new stands. The techs are getting into the swing of it and should make good progress next week.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 9/15/2003   email Fermilab